Below is a lift at Gugulethu mall depicted. The lift travels from the ground floor to the
eighth floor as shown. You have made studies during its installation and found out that
the cable has a mass of 6.3 kg/m. The lift cage has a mass of 750 kg and there are
three people in the lift, two weighing 75 kg each. The third person is you. Answer the
following questions without neglecting the weight of the third person in the lift
1
Expert's answer
2020-12-14T07:19:37-0500

I think the picture must be given but I try to solve

The lift travel ground floor to eighth floor. I have made a studies during its installation

and I found that the mass of cable is that 6.3 kg each meter. the mass of lift cage 750kg

and there are three people one of them is me and my mass is neglected, other two people mass is 75kg each and total mass of them are 150kg.

The lift travel to eighth floor begging and we take the length of seven floor

I consider the lift direction is upward x > 0, the velocity is v > a and acceleration a > 0 = 5m/s2

the total weight of system P = (M + 2*m)*g + (M + 2*m) * a = (750 + 75 * 2) (10 + 5)= 13500N;